# LLM-driven Indoor Scene Layout Generation via Scaled Human-aligned Data Synthesis and Multi-Stage Preference Optimization

### v1

@@ 1 Introduction
Indoor Scene Layout Generation has attracted growing attention across the 3D vision fields, including virtual environment synthesis [34, 10, 39, 4, 3], robot navigation [26, 25, 41], digital twin construction [5, 17, 27], interior design automation [36, 12, 14], and human-centric simulation [40, 37, 32, 18]. With the increasing need for simulating realistic virtual environments, learning to generate functionally plausible, physically feasible, and semantically meaningful spatial arrangements has become an essential capability for intelligent systems.
@@ 3 Methodology: 3D-SynthPlace & OptiScene
In this section, we present the 3D-SynthPlace dataset and OptiScene framework (as Fig. 2). We begin by providing a concise formulation of our layout generation problem in section 3.1. In section 3.2, we describe the dataset construction process and highlight the criteria and techniques used to ensure high-quality and diverse scene representations. Following this, in section 3.3, we detail the supervised fine-tuning (SFT) process, including the model’s input/output format and the design of the meta instruction prompt. And finally in section 3.4, we discuss how to use the Direct Preference Optimization (DPO) to do the iterative optimization to make the SFT scene layout result align with the human preferences.
@@ 4.2 Experiment results
Usability consistency. To evaluate the robustness and reliability of different layout generation methods, we conduct a usability consistency test in which each model generates 50 layouts per room type, and we record the proportion of layouts deemed valid (i.e., free of collisions, interpenetrations, and major structural violations). As shown in Tab. 3, OptiScene significantly outperforms all baselines in both bedroom and living room categories, achieving 75% and 50% usable layouts respectively. In contrast, DiffuScene and LLplace achieve much lower success rates, especially on complex scenes such as kitchens and bathrooms. While Holodeck performs reasonably well in bedrooms, its outputs remain less stable across room types.
Qualitative analysis. Fig. 4 shows visual comparisons across models on six room categories from 3D-Front and 3D-SynthPlace. I-Design consistently fails to construct scenes with more than 5–6 objects, often missing essential furniture (marked in blue), highlighting its limited scalability. DiffuScene and LLplace trained only on 3D-Front struggle to generalize, especially to novel room types like kitchens and bathrooms. DiffuScene in particular fails entirely in these categories. With 3D-SynthPlace training, both methods see noticeable improvement, yet issues like collisions and misalignments (red boxes) persist. In contrast, our proposed OptiScene generates coherent, well-structured layouts across all scenes, with correct functional grouping and minimal physical violations, demonstrating superior generalization and human-aligned spatial reasoning.
@@ 4.3 Downstream tasks
Room editing is essential for interactive layout generation, yet prior methods like DiffuScene and I-Design lack such capability. Inspired by LLplace, we extend 3D-SynthPlace with an object-level editing dataset containing addition and removal instructions. We fine-tune the model with mixed editing and original data, and evaluate on 20 edit cases using object-overlap rate (OOR) and editing success rate (ESR). As shown in Fig. 5, OptiScene maintains low OOR post-editing and achieves 90% ESR, with only two failed cases.
@@
Object-Centric navigation evaluates whether the generated layouts support goal-directed robot exploration. Given a command like “Find a <object> in the room,” the robot succeeds if the target appears within ±30° of its field of view and within 2 meters, measured by success rate (SR). We also report navigation error (NE), defined as the final distance to the target. As shown in Fig. 5, our layouts support stable navigation across LLM backends, indicating strong spatial usability.
@@ 4.4 Ablation study
We investigate whether prompting the model to generate high-level spatial semantics before layout prediction improves overall quality. As shown in Tab. 6, adding this reasoning step to the SFT model increases layout usability (from 28% to 33%). This demonstrates that explicitly guiding the model to first consider object relationships and room intent leads to more coherent and human-aligned scene structures.
@@
As shown in Tab. 6, adding reasoning improves OOR and UR by encouraging coherent object planning. DPO-Stage1 aligns outputs with human preferences, further reducing violations, while DPO-Stage2 introduces targeted noise to correct fine-grained physical errors. Qualitative results in Fig. 5 illustrate progressive improvement in layout realism and usability across stages.
@@
We compare two reasoning supervision strategies: high-level semantic descriptions and scene graphs. As shown in Fig. 7, using natural language descriptions leads to significantly better layout quality, with more coherent and functional arrangements. This is because LLMs are inherently better at interpreting and generating natural language than symbolic graph structures. Scene graphs, while explicit, lack global spatial context and are less aligned with the model’s pretraining distribution, resulting in fragmented or rigid layouts.
@@ A.2 The deleted Holodeck generated data
We also illustrate some deleted layouts to show the error of the room scenes in Fig. 8. In addition to filtering out invalid layouts, we exclude small objects produced during the Holodeck generation process, as they are not compatible with the problem definition of our room layout formulation.
@@ A.3 Data distribution description
The Fig. 9 presents a comprehensive distribution comparison between the 3D-Front and 3D-SynthPlace datasets across various dimensions:
Number of Objects Fig. 9(a): The box plot illustrates the distribution of furniture counts in different room types for both datasets. The 3D-Front dataset shows a higher median count in living rooms compared to bedrooms, while the 3D-SynthPlace dataset exhibits a similar pattern with slightly higher counts in living rooms and kitchens.
Room Distribution Fig. 9(b & c): Pie charts depict the room distribution within each dataset. The 3D-Front dataset is predominantly composed of living rooms (66.1%), followed by bedrooms (33.9%). In contrast, the 3D-SynthPlace dataset has a more diverse distribution, with living rooms (38%) and bedrooms (27.5%) being the most common, alongside significant portions of bathrooms (17.5%) and kitchens (17.1%).
Comparison of Number of Assets Fig. 9(d): A bar chart compares the total number of assets across different room types. The 3D-Front dataset has a significantly higher total number of assets, especially in living rooms, compared to the Holodeck-Synth dataset.
Comparison of Number of Classes Fig. 9(e): Another bar chart compares the number of classes available in each dataset. The 3D-Front dataset generally has more classes across different room types, with the most notable difference in the living room category.
@@ A.4 The reasoning description illustration
In Fig. 10, we show some reasoning high-level semantic descriptions with the 3D scene layouts. Examples of generated room layouts paired with natural language descriptions. Each row shows two 3D layout visualizations of the same room from different angles, accompanied by a GPT-generated description of the spatial arrangement and object relations. The descriptions emphasize functionality, object alignment, and spatial reasoning without relying on directional terms like "left" or "right".
@@ Appendix C More details of DPO
In Section 3.4, we introduce a two-stage DPO (Direct Preference Optimization) framework designed to align generated room layouts with human preferences. We begin by selecting 1,100 scene layouts from the Holodeck-Synth dataset and another 1,100 from the 3D-Front dataset. These layouts are characterized by reasonable, functional, and human-aligned spatial arrangements. In the second stage of DPO training (DPO-stage2), we generate hard negative samples by adding targeted spatial violations to the positive examples. Specifically, as illustrated in the Fig. 11, we introduce two types of noise: (1) Out-of-boundary violations, where furniture items extend beyond the floor area; (2) Object-overlap violations, where furniture pieces intersect or collide with each other. Despite improvements from SFT and DPO-stage1, the model still occasionally produces such errors. Therefore, we further fine-tune the model using these challenging negative samples to enhance spatial reasoning and layout robustness.
@@ D.2 Details of the evaluation metrics
Fig. 12 and Fig. 13 present additional qualitative comparisons between the outputs of our OptiScene model (left two columns) and the ground truth (right two columns) across various room configurations. It is important to emphasize that the objective of OptiScene is not to converge to the ground truth layouts exactly, but rather to generate plausible and functional indoor scenes based on high-level semantic and spatial constraints. The generated layouts exhibit a even more efficient use of space, better object alignment, and clearer functional zoning compared to the ground truth. These examples demonstrate the model’s capacity to generalize and produce reasonable, sometimes even preferable, alternative layouts that remain faithful to the intended room semantics. This highlights the potential of OptiScene to support diverse and creative room layout configurations.
@@
In Fig. 14 and Fig. 15, we illustrate more downstream tasks results with room editing and robot navigation. As shown in figures, with appropriate instruction, the model has the good capability to edit the room layout and generate the robot navigation path.

### v2

@@ 1 Introduction
Indoor Scene Layout Generation has attracted growing attention across the 3D vision fields, including virtual environment synthesis [36, 10, 42, 4, 3], robot navigation [26, 25, 44, 41], digital twin construction [5, 17, 27, 35], interior design automation [38, 12, 14], and human-centric simulation [43, 39, 32, 18, 34]. With the increasing need for simulating realistic virtual environments, learning to generate functionally plausible, physically feasible, and semantically meaningful spatial arrangements has become an essential capability for intelligent systems.
@@ 3 Methodology: 3D-SynthPlace & OptiScene
In this section, we present the 3D-SynthPlace dataset and OptiScene framework (as Fig.˜2). We begin by providing a concise formulation of our layout generation problem in section 3.1. In section 3.2, we describe the dataset construction process and highlight the criteria and techniques used to ensure high-quality and diverse scene representations. Following this, in section 3.3, we detail the supervised fine-tuning (SFT) process, including the model’s input/output format and the design of the meta instruction prompt. And finally in section 3.4, we discuss how to use the Direct Preference Optimization (DPO) to do the iterative optimization to make the SFT scene layout result align with the human preferences.
@@ 4.2 Experiment results
Usability consistency. To evaluate the robustness and reliability of different layout generation methods, we conduct a usability consistency test in which each model generates 50 layouts per room type, and we record the proportion of layouts deemed valid (i.e., free of collisions, interpenetrations, and major structural violations). As shown in Tab.˜3, OptiScene significantly outperforms all baselines in both bedroom and living room categories, achieving 75% and 50% usable layouts respectively. In contrast, DiffuScene and LLplace achieve much lower success rates, especially on complex scenes such as kitchens and bathrooms. While Holodeck performs reasonably well in bedrooms, its outputs remain less stable across room types.
Qualitative analysis. Fig.˜4 shows visual comparisons across models on six room categories from 3D-Front and 3D-SynthPlace. I-Design consistently fails to construct scenes with more than 5–6 objects, often missing essential furniture (marked in blue), highlighting its limited scalability. DiffuScene and LLplace trained only on 3D-Front struggle to generalize, especially to novel room types like kitchens and bathrooms. DiffuScene in particular fails entirely in these categories. With 3D-SynthPlace training, both methods see noticeable improvement, yet issues like collisions and misalignments (red boxes) persist. In contrast, our proposed OptiScene generates coherent, well-structured layouts across all scenes, with correct functional grouping and minimal physical violations, demonstrating superior generalization and human-aligned spatial reasoning.
@@ 4.3 Downstream tasks
Room editing is essential for interactive layout generation, yet prior methods like DiffuScene and I-Design lack such capability. Inspired by LLplace, we extend 3D-SynthPlace with an object-level editing dataset containing addition and removal instructions. We fine-tune the model with mixed editing and original data, and evaluate on 20 edit cases using object-overlap rate (OOR) and editing success rate (ESR). As shown in Fig.˜5, OptiScene maintains low OOR post-editing and achieves 90% ESR, with only two failed cases.
@@
Object-Centric navigation evaluates whether the generated layouts support goal-directed robot exploration. Given a command like “Find a <object> in the room,” the robot succeeds if the target appears within ±30° of its field of view and within 2 meters, measured by success rate (SR). We also report navigation error (NE), defined as the final distance to the target. As shown in Fig.˜5, our layouts support stable navigation across LLM backends, indicating strong spatial usability.
@@ 4.4 Ablation study
We investigate whether prompting the model to generate high-level spatial semantics before layout prediction improves overall quality. As shown in Tab.˜6, adding this reasoning step to the SFT model increases layout usability (from 28% to 33%). This demonstrates that explicitly guiding the model to first consider object relationships and room intent leads to more coherent and human-aligned scene structures.
@@
As shown in Tab.˜6, adding reasoning improves OOR and UR by encouraging coherent object planning. DPO-Stage1 aligns outputs with human preferences, further reducing violations, while DPO-Stage2 introduces targeted noise to correct fine-grained physical errors. Qualitative results in Fig.˜5 illustrate progressive improvement in layout realism and usability across stages.
@@
We compare two reasoning supervision strategies: high-level semantic descriptions and scene graphs. As shown in Fig.˜7, using natural language descriptions leads to significantly better layout quality, with more coherent and functional arrangements. This is because LLMs are inherently better at interpreting and generating natural language than symbolic graph structures. Scene graphs, while explicit, lack global spatial context and are less aligned with the model’s pretraining distribution, resulting in fragmented or rigid layouts.
@@ A.2 The deleted Holodeck generated data
We also illustrate some deleted layouts to show the error of the room scenes in Fig.˜8. In addition to filtering out invalid layouts, we exclude small objects produced during the Holodeck generation process, as they are not compatible with the problem definition of our room layout formulation.
@@ A.3 Data distribution description
The Fig.˜9 presents a comprehensive distribution comparison between the 3D-Front and 3D-SynthPlace datasets across various dimensions:
Number of Objects Fig.˜9(a): The box plot illustrates the distribution of furniture counts in different room types for both datasets. The 3D-Front dataset shows a higher median count in living rooms compared to bedrooms, while the 3D-SynthPlace dataset exhibits a similar pattern with slightly higher counts in living rooms and kitchens.
Room Distribution Fig.˜9(b & c): Pie charts depict the room distribution within each dataset. The 3D-Front dataset is predominantly composed of living rooms (66.1%), followed by bedrooms (33.9%). In contrast, the 3D-SynthPlace dataset has a more diverse distribution, with living rooms (38%) and bedrooms (27.5%) being the most common, alongside significant portions of bathrooms (17.5%) and kitchens (17.1%).
Comparison of Number of Assets Fig.˜9(d): A bar chart compares the total number of assets across different room types. The 3D-Front dataset has a significantly higher total number of assets, especially in living rooms, compared to the Holodeck-Synth dataset.
Comparison of Number of Classes Fig.˜9(e): Another bar chart compares the number of classes available in each dataset. The 3D-Front dataset generally has more classes across different room types, with the most notable difference in the living room category.
@@ A.4 The reasoning description illustration
In Fig.˜10, we show some reasoning high-level semantic descriptions with the 3D scene layouts. Examples of generated room layouts paired with natural language descriptions. Each row shows two 3D layout visualizations of the same room from different angles, accompanied by a GPT-generated description of the spatial arrangement and object relations. The descriptions emphasize functionality, object alignment, and spatial reasoning without relying on directional terms like "left" or "right".
@@ Appendix C More details of DPO
In Section 3.4, we introduce a two-stage DPO (Direct Preference Optimization) framework designed to align generated room layouts with human preferences. We begin by selecting 1,100 scene layouts from the Holodeck-Synth dataset and another 1,100 from the 3D-Front dataset. These layouts are characterized by reasonable, functional, and human-aligned spatial arrangements. In the second stage of DPO training (DPO-stage2), we generate hard negative samples by adding targeted spatial violations to the positive examples. Specifically, as illustrated in the Fig.˜11, we introduce two types of noise: (1) Out-of-boundary violations, where furniture items extend beyond the floor area; (2) Object-overlap violations, where furniture pieces intersect or collide with each other. Despite improvements from SFT and DPO-stage1, the model still occasionally produces such errors. Therefore, we further fine-tune the model using these challenging negative samples to enhance spatial reasoning and layout robustness.
@@ D.2 Details of the evaluation metrics
Fig.˜12 and Fig.˜13 present additional qualitative comparisons between the outputs of our OptiScene model (left two columns) and the ground truth (right two columns) across various room configurations. It is important to emphasize that the objective of OptiScene is not to converge to the ground truth layouts exactly, but rather to generate plausible and functional indoor scenes based on high-level semantic and spatial constraints. The generated layouts exhibit a even more efficient use of space, better object alignment, and clearer functional zoning compared to the ground truth. These examples demonstrate the model’s capacity to generalize and produce reasonable, sometimes even preferable, alternative layouts that remain faithful to the intended room semantics. This highlights the potential of OptiScene to support diverse and creative room layout configurations.
@@
In Fig.˜14 and Fig.˜15, we illustrate more downstream tasks results with room editing and robot navigation. As shown in figures, with appropriate instruction, the model has the good capability to edit the room layout and generate the robot navigation path.